Y=3x-4 (-3, -5) determine whether the ordered pair is a solution of the given equation. Please help me with this question
Answer by funmath(2925)

In order to determine if this ordered pair is a solution, substitute the appropriate values in for x and y into the equation and see if both sides equal. If the do, it's a solution. If not, it's not.
In this case x=-3 and y=-5
The y=3x-4 becomes
-5=3(-3)-4
-5=-9-4
-5=-13
Both sides are not equal, therefore (-3,-5) is not a solution of the given equation.